a room is 4 feet longer than its wide. its area is 192 square feet. what is the length of the room?
Answer by Alan3354(69443)

Try pairs of factors of 192, look for a pair that differ by 4.
------
Or,
Area = W*L = 192
L = W + 4
W*(W + 4) = 192
W^2 + 4W - 192 = 0
----
Now try pairs of factors of 192, look for a pair that differ by 4.